### v1

@@ 4.4.2 Cosmological correlators
Similar singularity analysis can also be applied to cosmological correlators. Let us consider an example of a triangle diagram (also known as the one-loop three-site graph) for the conformally coupled scalar in Friedmann–Lemaître–Robertson–Walker cosmology. Following the notation of [84], the corresponding correlator expressed in Baikov representation is proportional to
@@
and record the rest in SOFIA_ examples.nb [3].
To our knowledge, the singularity set for the triangle integral was never obtained before. Partial results exist [85] and we have checked that our singularity set contains them.
@@ 5 Future directions
This first implementation already pushes beyond the current state of the art in many cases. In particular, the ancillary file [3] includes several new predictions, including those of Fig. 1. These results highlight both the strength and versatility of [Uncaptioned image] , leading to results that were previously beyond reach.
@@
One of the most pressing questions is the algorithmic determination of the nature of singularities, i.e., the exponents $A$ and $B$ in the local behavior $f_{j}^{A}\log^{B}f_{j}$ as $f_{j}\to 0$. These exponents depend on the spacetime dimension and the geometry of the pinch manifold. In the case where the masses are sufficiently generic, this geometry is simple to describe, and consequently a closed-form formula for the exponents $A$ and $B$ is known [9]; see, e.g., [44, 86] for reviews. In “non-generic” situations, which are often the ones of physical interest, more work has to be done in analyzing the pinch manifold. Performing this analysis algorithmically would provide a significant improvement over the capabilities of [Uncaptioned image] , especially given that it would allow one to distinguish between square-root and logarithmic letters, thus optimizing the size of the proposed symbol alphabet.
@@
A related question concerns the Riemann sheet structure of Feynman integrals and scattering amplitudes. The simplest version of this question is to determine which singularities lie on the physical sheet or in the physical region. This question is usually addressed by either determining whether the loop momenta along the pinch manifold are physical, or whether the associated Schwinger parameters are non-negative. We have not attempted to address these questions since the Fubini reduction algorithm circumvents the need to solve the Landau equations directly. Therefore, it does not directly reconstruct the pinch manifold. Similarly, the discontinuity (monodromy) structure of Feynman integrals is more difficult to study in Baikov-like representations, as has been recognized for a long time [38, 87], since the integration contour itself depends on the kinematics. One possible direction is to keep track of the genealogy of singularities in the reduction algorithm, see [32, Sec. 6], which is closely related to the hierarchical structure; see, e.g., [88].
The questions about the sheet structure have their avatar in the symbol properties for polylogarithmic integrals. For example, presence on the physical sheet corresponds to the first-entry condition and the discontinuity structure is encoded in symbol adjacency. See [89, 90] for recent work on the connection to Landau analysis. We believe that the same information is encoded in the commutation relations of the matrices $\mathbf{\Omega}_{i}$ in (1.1) and boundary conditions even for non-polylogarithmic integrals, which means being able to predict it independently would put constraints on the $\mathbf{\Omega}_{i}$’s. See [70, 91] for explicit examples in the context of extended Steinmann relations.
@@
To find the symbol alphabet we wrapped in [Uncaptioned image] the publicly available package Effortless [36]. More work is needed in the algorithmic determination of alphabets, which is currently the main bottleneck in [Uncaptioned image] as can be seen in the timings indicated in Fig. 1. We have already mentioned that this procedure would benefit from determining algorithmically which letters are square-root vs. logarithmic, though this information can be provided through the option AddLetters, as explained in Sec. 3.3.1. In addition, more conceptual work is needed to understand how general letters can be, including the maximum multiplicity of square roots that can be expected and which singularities are allowed to appear as root arguments. Here, we committed to the ansatz (2.36) with double square roots. More general ansätze can be implemented as needed.
@@
Specializing the spacetime dimension may prove useful in other contexts, such as the S-matrix bootstrap [92], where understanding the singularity structure can be essential to make use of tools such as dispersion relations, particularly at higher points [13, 93] and beyond lightest-particle scattering [94, 95]. Another relevant application is the study of two-dimensional integrable models, in which anomalous thresholds manifest as Coleman–Thun poles [96, 97]. For example, an automated tool like [Uncaptioned image] could help extend recent studies on perturbative integrability beyond one-loop [98, 99].
@@
One main drawback of the loop-by-loop Baikov approach in these cases is that the resulting $N$ in (2.6) typically remains high, because each $\mathcal{E}_{a}$ is as large as possible for maximally connected diagrams, regardless of our minimizing strategy for the choice of LoopEdges. This poses a computational challenge starting at three-loops, because maximally connected diagrams lead to systems with a large number ($\geqslant 5$) of variables to eliminate, even on the maximal cut. An example (with $N=5$ on the maximal cut) is the symmetric “envelope diagram” from [66, Fig. 1e], whose singularities were previously calculated with PLD.jl in Schwinger parameter space for different masses assignments. Although the symbolic extraction of singularities for this diagram currently requires an impractical runtime in [Uncaptioned image] , we were able to use it to successfully reproduce numerically all of its known singularities.
@@
Finally, because the connectivity requirement naturally forces a uniformity in the distribution of vertices, maximally connected diagrams are generally highly symmetric. Fully exploiting this property, as demonstrated in [34, 35] for the physical sheet singularity, could plausibly lead to further improvements of [Uncaptioned image] ’s performance on these graphs.
@@ Appendix A Derivation of Baikov representations
This is because there are at most $n-1$ independent $p_{i}$’s due to momentum conservation. In addition, the number of independent $p_{i}$’s cannot exceed the spacetime dimension $\mathrm{D}_{0}$ (in dimensional regularization, $\mathrm{D}=\mathrm{D}_{0}-2\varepsilon$). The total number of such independent momenta $q_{i}$ is therefore $M=L+\mathcal{E}$. They define a natural coordinate system, where we work in terms of the Lorentz-invariant scalar products: $X_{ij}=q_{i}\cdot q_{j}$.
@@
$\begin{split}\mathrm{d}^{M-a}\ell_{a\parallel}&=\sqrt{\det G}\prod_{b=1}^{M-a}\mathrm{d}\lambda_{ab}\,,\end{split}$ (A.12)
@@
$\int\limits_{S_{\mathrm{D}-M+a}}\mathrm{d}^{\mathrm{D}-M+a}\ell_{a\perp}=\left[\frac{\det G(\ell_{a},\ell_{a+1},\dots,p_{\mathcal{E}})}{\det G(\ell_{a+1},\dots,p_{\mathcal{E}})}\right]^{\frac{\mathrm{D}-M+a-2}{2}}\frac{\Omega_{\mathrm{D}-M+a}\,\mathrm{d}X_{aa}}{2}\,,$ (A.18)
@@
$\begin{split}\int\mathrm{d}^{\mathrm{D}}\ell_{1}\mathrm{d}^{\mathrm{D}}\ell_{2}\dots\mathrm{d}^{\mathrm{D}}&\ell_{L}\,\mathcal{I}(\ell_{1},\dots,\ell_{L})=\frac{\pi^{\frac{L(1+2(\mathrm{D}+\mathcal{E})-L)}{4}}}{\prod_{l=1}^{L}\Gamma\left(\frac{\mathrm{D}-M+l}{2}\right)}\det G(p_{1},\dots,p_{\mathcal{E}})^{(-\mathrm{D}+\mathcal{E}+1)/2}\\ &\times\int_{\Gamma}\prod_{i=1}^{L}\prod_{j=i}^{M}\mathrm{d}X_{ij}\det G(\ell_{1},\dots,\ell_{L},p_{1},\dots,p_{\mathcal{E}})^{(\mathrm{D}-M-1)/2}f(X_{ij})\,,\end{split}$ (A.21)
where the multiple factors of $i\pi^{\mathrm{D}/2}$ in (2.1) were absorbed in the integrand $f$. Moreover, the integration domain $\Gamma$ is given by imposing $L$ conditions given in (A.20), such that $\Gamma\equiv\Gamma_{1}\cap\Gamma_{2}\cap\cdots\cap\Gamma_{L}=\{\det G(\ell_{1},\dots,\ell_{L},p_{1},\dots,p_{\mathcal{E}})>0\}$. [^2] The last equality follows from the observation that, since $\det G(q_{1},q_{2},\dots,q_{n})>0$, the biggest Gram matrix is positive definite. A key property of positive definite matrices is that every principal submatrix is also positive definite. In particular, Sylvester’s criterion tells us that all leading principal minors are positive, and in fact, this extends to any principal submatrix. Therefore, the submatrix corresponding to $q_{2},q_{3},\dots,q_{n}$ is positive definite, ensuring that $\det G(q_{2},q_{3},\dots,q_{n})>0$ and so on. In contrast, in the loop-by-loop case discussed in the main text, the multiple pairs of internal-external Gram matrices typically are not principal submatrices of one another. This occurs when they satisfy (A.21), but with Gram matrices of different sizes (see, e.g., (2.15)). In such cases, the integration contour remains the complete intersection of the corresponding $\Gamma_{i}$ regions.
@@ Appendix B Interface with Julia
Running this command successfully for the first time requires a few steps. First, the user must ensure that PLD.jl and all its dependencies are correctly installed and running on their Julia interface. The installation procedure can be found in [65].
@@ Appendix C Subtleties with kinematic replacements
To illustrate this point, we consider the two acnode diagrams defined by:
@@
For all the details of this example, see the notebook SOFIA_ examples.nb in [3]. The $\delta$-procedure outlined above is implemented and executed automatically by [Uncaptioned image] .

### v2

@@ 4.4.2 Cosmological correlators
Similar singularity analysis can also be applied to cosmological correlators. Let us consider an example of a triangle diagram (also known as the one-loop three-site graph) for the conformally coupled scalar in Friedmann–Lemaître–Robertson–Walker cosmology. Following the notation of [Benincasa:2024lxe], the corresponding correlator expressed in Baikov representation is proportional to
@@
and record the rest in SOFIA_examples.nb [repo].
To our knowledge, the singularity set for the triangle integral was never obtained before. Partial results exist [Benincasa:2024ptf] and we have checked that our singularity set contains them.
@@ 5 Future directions
This first implementation already pushes beyond the current state of the art in many cases. In particular, the ancillary file [repo] includes several new predictions, including those of Fig. 1. These results highlight both the strength and versatility of [Uncaptioned image] , leading to results that were previously beyond reach.
@@
One of the most pressing questions is the algorithmic determination of the nature of singularities, i.e., the exponents $A$ and $B$ in the local behavior $f_{j}^{A}\log^{B}f_{j}$ as $f_{j}\to 0$. These exponents depend on the spacetime dimension and the geometry of the pinch manifold. In the case where the masses are sufficiently generic, this geometry is simple to describe, and consequently a closed-form formula for the exponents $A$ and $B$ is known [Landau:1959fi]; see, e.g., [pham2011singularities, Hannesdottir:2022bmo] for reviews. In “non-generic” situations, which are often the ones of physical interest, more work has to be done in analyzing the pinch manifold. Performing this analysis algorithmically would provide a significant improvement over the capabilities of [Uncaptioned image] , especially given that it would allow one to distinguish between square-root and logarithmic letters, thus optimizing the size of the proposed symbol alphabet.
@@
A related question concerns the Riemann sheet structure of Feynman integrals and scattering amplitudes. The simplest version of this question is to determine which singularities lie on the physical sheet or in the physical region. This question is usually addressed by either determining whether the loop momenta along the pinch manifold are physical, or whether the associated Schwinger parameters are non-negative. We have not attempted to address these questions since the Fubini reduction algorithm circumvents the need to solve the Landau equations directly. Therefore, it does not directly reconstruct the pinch manifold. Similarly, the discontinuity (monodromy) structure of Feynman integrals is more difficult to study in Baikov-like representations, as has been recognized for a long time [Gribov:1962ft, islam1965analytic] (see [Tourkine:2023xtu] for a recent study), since the integration contour itself depends on the kinematics. One possible direction is to keep track of the genealogy of singularities in the reduction algorithm, see [Brown:2009ta, Sec. 6], which is closely related to the hierarchical structure; see, e.g., [Berghoff:2022mqu].
The questions about the sheet structure have their avatar in the symbol properties for polylogarithmic integrals. For example, presence on the physical sheet corresponds to the first-entry condition and the discontinuity structure is encoded in symbol adjacency. See [Hannesdottir:2024cnn, Hannesdottir:2024hke] for recent work on the connection to Landau analysis. We believe that the same information is encoded in the commutation relations of the matrices $\mathbf{\Omega}_{i}$ in (1.1) and boundary conditions even for non-polylogarithmic integrals, which means being able to predict it independently would put constraints on the $\mathbf{\Omega}_{i}$’s. See [Abreu:2020jxa, Abreu:2021smk] for explicit examples in the context of extended Steinmann relations.
@@
To find the symbol alphabet we wrapped in [Uncaptioned image] the publicly available package Effortless [Effortlessxxx]. More work is needed in the algorithmic determination of alphabets, which is currently the main bottleneck in [Uncaptioned image] as can be seen in the timings indicated in Fig. 1. We have already mentioned that this procedure would benefit from determining algorithmically which letters are square-root vs. logarithmic, though this information can be provided through the option AddLetters, as explained in Sec. 3.3.1. In addition, more conceptual work is needed to understand how general letters can be, including the maximum multiplicity of square roots that can be expected and which singularities are allowed to appear as root arguments. Here, we committed to the ansatz (2.36) with double square roots. More general ansätze can be implemented as needed.
@@
Specializing the spacetime dimension may prove useful in other contexts, such as the S-matrix bootstrap [Kruczenski:2022lot], where understanding the singularity structure can be essential to make use of tools such as dispersion relations, particularly at higher points [Caron-Huot:2023ikn, Guerrieri:2024ckc] and beyond lightest-particle scattering [Homrich:2019cbt, Correia:2022dcu]. See also [Vergu:2023rqz, Bargiela:2024hgt] for recent progress on computation of Feynman integrals via dispersion relations. Another relevant application is the study of two-dimensional integrable models, in which anomalous thresholds manifest as Coleman–Thun poles [Coleman:1978kk, Braden:1990wx]. For example, an automated tool like [Uncaptioned image] could help extend recent studies on perturbative integrability beyond one-loop [Dorey:2022fvs, Fabri:2024xmi].
@@
One main drawback of the loop-by-loop Baikov approach in these cases is that the resulting $N$ in (2.6) typically remains high, because each $\mathcal{E}_{a}$ is as large as possible for maximally connected diagrams, regardless of our minimizing strategy for the choice of LoopEdges. This poses a computational challenge starting at three-loops, because maximally connected diagrams lead to systems with a large number ($\geqslant 5$) of variables to eliminate, even on the maximal cut. An example (with $N=5$ on the maximal cut) is the symmetric “envelope diagram” from [Mizera:2021icv, Fig. 1e], whose singularities were previously calculated with PLD.jl in Schwinger parameter space for different masses assignments. Although the symbolic extraction of singularities for this diagram currently requires an impractical runtime in [Uncaptioned image] , we were able to use it to successfully reproduce numerically all of its known singularities.
@@
Finally, because the connectivity requirement naturally forces a uniformity in the distribution of vertices, maximally connected diagrams are generally highly symmetric. Fully exploiting this property, as demonstrated in [KORS, Correia:2021etg] for the physical sheet singularity, could plausibly lead to further improvements of [Uncaptioned image] ’s performance on these graphs.
@@ Appendix A Derivation of Baikov representations
This is because there are at most n1n-1 independent $p_{i}$’s due to momentum conservation. In addition, the number of independent $p_{i}$’s cannot exceed the spacetime dimension $\mathrm{D}_{0}$ (in dimensional regularization, $\mathrm{D}=\mathrm{D}_{0}-2\varepsilon$). The total number of such independent momenta $q_{i}$ is therefore $M=L+\mathcal{E}$. They define a natural coordinate system, where we work in terms of the Lorentz-invariant scalar products: $X_{ij}=q_{i}\cdot q_{j}$.
@@
[…]
@@
$\hskip-5.69046pt\int\limits_{S_{\mathrm{D}-M+a}}\hskip-8.5359pt\mathrm{d}^{\mathrm{D}-M+a}\ell_{a\perp}=\left[\frac{\det G(\ell_{a},\ell_{a+1},\dots,p_{\mathcal{E}})}{\det G(\ell_{a+1},\dots,p_{\mathcal{E}})}\right]^{\frac{\mathrm{D}-M+a-2}{2}}\frac{\Omega_{\mathrm{D}-M+a}\,\mathrm{d}X_{aa}}{2}\,$, (A.18)
@@
[…]
where the multiple factors of $i\pi^{\mathrm{D}/2}$ in (2.1) were absorbed in the integrand $f$. Moreover, the integration domain $\Gamma$ is given by imposing $L$ conditions given in (A.20), such that $\Gamma\equiv\Gamma_{1}\cap\Gamma_{2}\cap\cdots\cap\Gamma_{L}=\{\det G(\ell_{1},\dots,\ell_{L},p_{1},\dots,p_{\mathcal{E}})>0\}$. [^2] The last equality follows from the observation that, since $\det G(q_{1},q_{2},\dots,q_{n})>0$, the biggest Gram matrix is positive definite. A key property of positive definite matrices is that every principal submatrix is also positive definite. In particular, Sylvester’s criterion [Horn_Johnson_1985] tells us that all leading principal minors are positive, and in fact, this extends to any principal submatrix. Therefore, the submatrix corresponding to $q_{2},q_{3},\dots,q_{n}$ is positive definite, ensuring that $\det G(q_{2},q_{3},\dots,q_{n})>0$ and so on. In contrast, in the loop-by-loop case discussed in the main text, the multiple pairs of internal-external Gram matrices typically are not principal submatrices of one another. This occurs when they satisfy (A.21), but with Gram matrices of different sizes (see, e.g., (2.15)). In such cases, the integration contour remains the complete intersection of the corresponding $\Gamma_{i}$ regions.
@@ Appendix B Interface with Julia
Running this command successfully for the first time requires a few steps. First, the user must ensure that PLD.jl and all its dependencies are correctly installed and running on their Julia interface. The installation procedure can be found in [MathRepoTutorial].
@@ Appendix C Subtleties with kinematic replacements
To illustrate this point, we consider the two acnode diagrams [Eden:1966dnq] defined by:
@@
For all the details of this example, see the notebook SOFIA_examples.nb in [repo]. The $\delta$-procedure outlined above is implemented and executed automatically by [Uncaptioned image] .